The difference between the digits of a digit number is 4. The sum of the numbers and the number formed by reversing its digit is 110. Find the two-digit number.
Answer by ankor@dixie-net.com(22740)

let a = the 10's digit
let b = the units
then
10a+b = the two-digit number
:
The difference between the digits of a digit number is 4.
a - b = 4
The sum of the numbers and the number formed by reversing its digit is 110.
10a+b + 10b+a = 110
11a + 11b = 110
simplify, divide by 11
a + b = 10
:
Use elimination on these two equations
a - b = 4
a + b = 10
-------------addition eliminates b, find a
2a = 14
a = 7
then
b = 7 - 4
b = 3
:
Find the two-digit number. 73 is the number
;
;
Check this: 73 + 37 = 110